Two cars are 15 kms apart. One is turning at a speed of 50kmph and the other at 40kmph . How much time will it take for the two cars to meet?
Found 2 solutions by mananth, richwmiller:
Answer by mananth(16949)

Two cars are 15 kms apart. One is turning at a speed of 50kmph and the other at 40kmph . How much time will it take for the two cars to meet?
The distance between them is reducing at the rate of 10 mph. (50-40)
so to meet it will take 1.5 hrs.

Answer by richwmiller(17219)

It is not clear if they are heading towards each other or heading to the same place and who is further ahead.
If they are heading towards each other.
50t+40t=15
90t=15
t=1/6 hr =10 minutes if heading towards each other

If the slower car is 15 km ahead
then 40t+15=50t
15=50t-40t
15=10t
15/10=t
1.5 hrs forthe 50 kph to catch up
if the 50 kph is ahead by 15k
50t+15=40t
15=-10t
15/-10=t
-1.5 hr=t
time can't go backwards so the slower car can never catch up.